OpenStudy (anonymous):

how to prove the following integral i think i know the way but i need some guidance through it please Integral x(x+a)^n dx = (x+a)^n+1 ((n+1) x - a) / (n+1)(n+2)

OpenStudy (anonymous):

this is how i tried to solve it let u=x u'=1dx v'=(x+a)^n v=integral (x+a)^n dx = (x+a)^n+1 / (n+1)

OpenStudy (anonymous):

am i right so far?

OpenStudy (sidsiddhartha):

yeah do u know integration by parts?

OpenStudy (anonymous):

yeah could you please see if im getting it right please

OpenStudy (anonymous):

= uv - integral u'v

OpenStudy (anonymous):

Integral x(x+a)^n dx =x(x+a)^n+1 / (n+1) - integral 1 * (x+a)^n+1 / (n+1) dx

OpenStudy (sidsiddhartha):

yeah but i think its not required here \[take~~u=x+a\\then ~~dx=du\] ok?

OpenStudy (anonymous):

ok

OpenStudy (sidsiddhartha):

and also x=u-a right

OpenStudy (anonymous):

could you please show it to me step by step

OpenStudy (anonymous):

ive got my notebook next to me

OpenStudy (sidsiddhartha):

yeah i'll show u

OpenStudy (anonymous):

thank you so much

OpenStudy (sidsiddhartha):

now try to replace them in the integral it will be-- \[\Large \int\limits_{}^{}x(x+a)^ndx=\int\limits_{}^{}(u-a)*u^ndu\] ok ?

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

what is u?

OpenStudy (sidsiddhartha):

\[\because ~(x+a)=u\\and ~~then ~~x=(u-a)\\and ~also ~(x+a)^n=u^n\\OK?\]

OpenStudy (anonymous):

understood

OpenStudy (sidsiddhartha):

just used u-substitution

OpenStudy (sidsiddhartha):

ok

OpenStudy (anonymous):

is there a trick to know this?

OpenStudy (sidsiddhartha):

nope u just have to think the possible ways to make it simpler now multiplying this we have\[\Large \int\limits_{}^{}[u^{n+1}-au^n]dx=\int\limits_{}^{}u^{n+1}dx-\int\limits_{}^{}au^ndx\] getting this?

OpenStudy (anonymous):

let me write it down

OpenStudy (sidsiddhartha):

sorry those dx will be du

OpenStudy (anonymous):

i didnt get this part

OpenStudy (sidsiddhartha):

\[\Large \int\limits_{}^{}u^{n+1}du-\int\limits_{}^{}au^n du\]

OpenStudy (anonymous):

ok i do now

OpenStudy (anonymous):

you multiplied

OpenStudy (sidsiddhartha):

yeah

OpenStudy (sidsiddhartha):

and from earlier assumption dx=du ok?

OpenStudy (sidsiddhartha):

now just simple integration

OpenStudy (sidsiddhartha):

using \[\Huge \int\limits_{}^{}x^ndx=\frac{ x^{n+1} }{ n+1 }+C\]

OpenStudy (anonymous):

how is dx = dy?

OpenStudy (anonymous):

ok i get it sorry

OpenStudy (anonymous):

because u = x+a u'= 1 dx

OpenStudy (sidsiddhartha):

look from the first step i took \[x+a=u\\now ~just~differentiating~both~sides~we~have\\dx=da+du\\but ~da=differentiation ~of~a~constant=0\\so ~just \\dx=du\]

OpenStudy (sidsiddhartha):

yeah thats correct :)

OpenStudy (anonymous):

what shall i do next ?

OpenStudy (anonymous):

add 1 to u^n+1

OpenStudy (anonymous):

/ by n+2

OpenStudy (anonymous):

i guess ill know how to do it from here

OpenStudy (sidsiddhartha):

yeah next use\[\int\limits_{}^{}x^{n}dx=\frac{ x^{n+1} }{ n+1 }\] so \[\large \int\limits_{}^{}u^{n+1}du=\frac{ u^{n+2} }{ n+2 }\] ok?

OpenStudy (anonymous):

does "a" go outside the integral?

OpenStudy (sidsiddhartha):

yeah it will \[\int\limits_{}^{}au^ndu=a \int\limits_{}^{}u^ndu=a*\frac{ u^{n+1} }{ n+1 }\]

OpenStudy (anonymous):

and then ill have to plug in u

OpenStudy (anonymous):

thank you so much for your time

OpenStudy (sidsiddhartha):

no problem buddy :)

OpenStudy (anonymous):

ill probably need help when putting it all together :)

OpenStudy (anonymous):

sorry for being a pain in the retrice:)

OpenStudy (sidsiddhartha):

lol ;)

OpenStudy (anonymous):

i took (n+1)(n+2) as the common denominator

OpenStudy (anonymous):

i meant retricehehehe

OpenStudy (anonymous):

i guess this website doesnt allow you to use such words lol

OpenStudy (anonymous):

i need help @sidsiddhartha

OpenStudy (anonymous):

@sidsiddhartha if you could please help me on the last bit

OpenStudy (sidsiddhartha):

yeah :)

OpenStudy (anonymous):

i solved it thanks :)

OpenStudy (anonymous):

that was proof number 7 of integral now im doing proof 8 which is integral 1 / (1+x^2) dx = inverse tan x

OpenStudy (anonymous):

i know that the differentiation of inverse tan x is 1 / (1+x^2)

OpenStudy (anonymous):

but how can we find the integral of 1 / (1+x^2) dx using substitution or any other method

OpenStudy (anonymous):

if we use u=1+x^2 then the derivative of u would have an x in it

OpenStudy (sidsiddhartha):

ok u have to use this form--\[\int\limits_{}^{}\frac{ dx }{ x^2+a^2 }=\frac{ 1 }{ a }\tan^{-1} \frac{ x }{ a}+c\]

OpenStudy (anonymous):

let me have a look at it for a moment

OpenStudy (anonymous):

ok so it is the same form because 1^2 is 1

OpenStudy (sidsiddhartha):

then--